This question isn't shown in the book.
Question:
The area of a square is 32 more than the perimeter. Find the length of a side.


If you could help, I would really appreciate it.
Thank You, Amy

Answer by checkley71(8403)

X^2=THE AREA
4X=PERIMETER THEN
X^2=4X+32
X^2-4X-32=0
(X-8)(X+4)=0
X-8=0
X=8 SOLUTION
PROOF
8*8=4*8+32
64=32+32
64=64